OpenStudy (anonymous):

Solve the equation 3x+8=23-3x

OpenStudy (anonymous):

6x=23-8 6x=15 x=15/6 x=5/2

OpenStudy (anonymous):

Could you break it down more?

OpenStudy (anonymous):

means??????

OpenStudy (anonymous):

Show your working

OpenStudy (anonymous):

adding 3x on both sides.... 6x+8=23 subtract 8 from both sides 6x=23-8 6x=15 x=15/6 x=5/2